# A new framework for nonlinear Kalman filters

### v10

@@ 3 Results
To demonstrate the effectiveness of the proposed framework, we validated it in five different applications suitable for nonlinear KFs. A brief description of the five applications is presented in Table 1, and a detailed description of the system models and parameter setup can be found in Appendix B. For each scenario, we combined the old and proposed framework with four different types of nonlinear KFs: EKF, EKF2, UKF, and CKF. We also implemented the IEKF for comparison. The detailed EKF, EKF2, UKF, and CKF algorithms used in the paper can be found in Algorithms 35 in Appendix C. All these nine $(4+4+1)$ nonlinear KFs are used for state estimation under different measurement noise setups (the covariance of the measurement noises differs in different setups). To guarantee the fairness of the comparison and reduce the effect of randomness, each KF is simulated 10,000 times for each measurement noise setup, and all the KFs use the same random seed during the simulation. This means that the random initial states’ errors, random process noises, and random measurement noises are the same for the $i^{th}$ simulation of EKF and the $i^{th}$ simulation of UKF. In other words, all these random noises were generated for the 10,000 runs a priori and recalled for simulating each estimation method.
@@
The root mean squared error (RMSE) of the state estimations given by different nonlinear KFs are shown in Figs. 37. In all figures, different types of nonlinear KFs are represented by different colors and markers, and the old and new frameworks are represented by dotted lines and solid lines, respectively. The x-axis represents the standard deviations of the measurements. The root mean squared error is calculated from the estimation errors in 10,000 simulations. The estimation error is calculated as the difference between the estimated state and the true state at the end of the final iteration of each simulation. Note that we only showed the RMSE of two representative states for systems with more than two states. For example, in 3D target tracking, the six states are the object’s speed and position along the three axes, and Fig. 3 examines the object’s speed and position estimation accuracy along the x-axis. As can be seen from the five figures, the simulation results are very similar to each other. Therefore, only the results of the first task, 3D target tracking, are analyzed in depth in this section.
@@
Figure 5: . The root mean squared error of the state estimations under different measurement noise setups (synchronous generator state estimation).
@@
By comparing the performance of IEKF and EKF (with and without the new framework) in Figs. 37, we can see that IEKF does perform better than EKF with the old framework, yet it is almost always worse than EKF with the new framework despite requiring longer computational time. As we mentioned in the Introduction, the main reason for such a difference is that IEKF still gives overconfident covariance estimation, especially when it fails to converge, which happens quite often in practice.
As we mentioned in the Introduction, the fundamental idea of the proposed framework is to make a more accurate covariance estimation by re-approximating the measurement functions after the state update. To validate if the covariance estimation really becomes more accurate in the new framework, we compared the actual and estimated RMSE of the x-axis position (the first state in the first system) in the old and new framework in Fig. 8. In the figure, for each type of nonlinear KF (represented by different line colors) and framework (the old and new frameworks are respectively shown in the top and bottom subfigures), the closeness between the actual RMSE (the solid line) and estimated RMSE (dotted line) represents the accuracy of the state covariance estimation. Apparently, when the standard deviations of the two measurements are lower than 0.1 meters, the conventional framework always significantly underestimates the covariance, sometimes by several orders of magnitude, regardless of the type of nonlinear KFs used. On the other hand, for the new framework, only the RMSE of EKF is underestimated, and the covariance estimation of other types of nonlinear KFs matches the actual value very well. This, again, can be explained by the fact that EKF neglects the second and higher-order information of the measurement functions, so its covariance estimation is less accurate than other methods. Nevertheless, the covariance estimation of EKF is still much more precise when combined with the new framework.
Figure 8: The actual and estimated root mean squared error of the x-axis position estimation.
Another interesting observation of Fig. 8 is that the estimated RMSE of the x-axis position (the dashed lines) is very similar for different combinations of frameworks and KFs. This suggests that the accuracy of the estimation mainly relates to whether the covariance estimation is accurate or not, which leads to the conclusion mentioned at the end of the Introduction — to achieve a more accurate estimation, finding the optimal Kalman gain is not as important as accurately estimating the covariance.
@@
Another important metric for evaluating the performance of different filters is the convergence characteristics. We visualized this metric by fixing the standard deviations of the two measurements to 0.01 m and plotted the RMSE of the state estimations after different numbers of iterations. The results are shown in Fig. 9. From the figure, we can see that all types of nonlinear KFs can converge faster after being combined with the new framework. This result suggests that the proposed framework is very suitable for applications that require high convergence speed.
@@
Since the extra steps introduced in the new framework take time, it is also important to compare the computational time of the two frameworks. The results are summarized in Fig. 10. Note that the average runtime shown in the figure is calculated from over 100,000 simulations run on the same computer. Specifically, for each combination of nonlinear KF and framework, we recorded the method’s runtime whenever it was used and calculated the average. After the average runtime is calculated, the value is normalized with respect to the average runtime of the conventional EKF for the same application. In other words, the normalized average runtime of “EKF (old)” always equals one in Fig. 10.
Figure 10: Comparison of the computational time of different methods and frameworks.
@@ 4 Conclusion
In this paper, we proposed a new framework that can be applied to all types of nonlinear KFs. The new framework re-approximates the measurement functions after the state update, solving the problem of overconfident state covariance estimation, which is an inherent problem in the conventional framework for nonlinear KFs. The new framework was tested in five different applications, reducing the estimation errors by more than an order of magnitude when the measurement noise is low and significantly reducing the convergence time of the algorithm. Considering that all types of sensors will become increasingly more accurate in the future, the proposed framework is expected to bring greater benefits for all kinds of nonlinear KFs over time.
@@ Appendix A Supplementary proof for Theorem 1
satisfy $\mathbb{E}[||x||^{2}]<\infty$ and $\mathbb{E}[||y||^{2}]<\infty$ (i.e., $x,y\in L^{2}(\Omega;\mathbb{R}^{n})$). Then, the operation on $L^{2}(\Omega;\mathbb{R}^{n})$ by
@@
Positive-definiteness is satisfied because for any $x\in L^{2}$, $\mathbb{E}[x^{T}x]\geq 0$, and the equality sign only holds when $x\equiv 0$.
@@
Let ${\color[rgb]{1,0,0}S}$ be a random symmetric and positive definite matrix whose mean is the identity matrix $I$. Let $\bm{v}$ be a random vector whose mean is $\bm{v}_{0}$. The following inequality holds:
$\mathbb{E}(\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v})\geq\bm{v}_{0}^{T}\bm{v}_{0}$ (22)
Note that the equality sign only holds when ${\color[rgb]{1,0,0}S^{-1}}\bm{v}$ is constant.
@@
Let $\bm{x}={\color[rgb]{1,0,0}S^{-\frac{1}{2}}}\bm{v},\bm{y}={\color[rgb]{1,0,0}S^{\frac{1}{2}}}\bm{v}_{0}$. From the definitions,
$\begin{cases}\mathbb{E}[\bm{x}^{T}\bm{x}]=\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v}]\\ \mathbb{E}[\bm{y}^{T}\bm{y}]=\mathbb{E}[\bm{v}_{0}^{T}{\color[rgb]{1,0,0}S}\bm{v}_{0}]=\bm{v}_{0}^{T}\bm{v}_{0}\\ \mathbb{E}[\bm{x}^{T}\bm{y}]=\mathbb{E}[\bm{v}^{T}\bm{v}_{0}]=\bm{v}_{0}^{T}\bm{v}_{0}\end{cases}$ (23)
According to Lemma 3, $\mathbb{E}[\bm{x}^{T}\bm{y}]$ is an inner product on $L^{2}(\Omega;\mathbb{R}^{n})$. Applying Cauchy–Schwarz inequality, we have:
$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v}]\geq\frac{(\bm{v}_{0}^{T}\bm{v}_{0})^{2}}{\bm{v}_{0}^{T}\bm{v}_{0}}$ (24)
which completes the proof. Note that the equality holds only when ${\color[rgb]{1,0,0}S^{-1}}\bm{v}$ is constant.
@@ Appendix B The detailed simulation setup
This example is adapted from Example 3.3.1 in [26]. As shown in Fig. 11(a), a target moves in 3D space, and two sensors are used to track its location. The first sensor is at the origin, and the second sensor moves in a circle. The second sensor’s location $(x_{s1},x_{s2},x_{s3})$ at step $k=1,2,\ldots,30$ is represented by:
$\begin{cases}x_{s1,k}=20+20\cos{\big{(}\frac{(k-1)\pi}{15}\big{)}}\\ x_{s2,k}=20+20\sin{\big{(}\frac{(k-1)\pi}{15}\big{)}}\\ x_{s3,k}=0\end{cases}$ (25)
The states are the target’s positions $(x_{1},x_{2},x_{3})$ and speeds $(v_{1},v_{2},v_{3})$ along the three axes. The measurements $y_{1},y_{2}$ are the distances between the target and the two sensors. The inputs are the target’s accelerations along the three axes, which are assumed to be zero in this example. When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (26,27). Note that the noises are omitted in the equations.
$\begin{bmatrix}x_{1,k}\\ x_{2,k}\\ x_{3,k}\\ v_{1,k}\\ v_{2,k}\\ v_{3,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}+v_{1,k-1}\\ x_{2,k-1}+v_{2,k-1}\\ x_{3,k-1}+v_{3,k-1}\\ v_{1,k-1}\\ v_{2,k-1}\\ v_{3,k-1}\end{bmatrix}$ (26)
$\begin{bmatrix}y_{1,k}\\ y_{2,k}\end{bmatrix}=\begin{bmatrix}\sqrt{\sum_{i=1}^{3}x_{i,k}^{2}}\\ \sqrt{\sum_{i=1}^{3}(x_{i,k}-x_{si,k})^{2}}\end{bmatrix}$ (27)

### v11

@@ 3 Results
To demonstrate the effectiveness of the proposed framework, we validate it in five different applications suitable for nonlinear KFs. A brief description of the five applications is presented in Table 1, and a detailed description of the system models and parameter setup can be found in Appendix B. For each scenario, we combine the old and proposed framework with four different types of nonlinear KFs: EKF, EKF2, UKF, and CKF. We also implemente the IEKF for comparison. The detailed EKF, IEKF, EKF2, UKF, and CKF algorithms used in the paper can be found in Algorithms 25 in Appendix C. All these nine $(4+4+1)$ nonlinear KFs are used for state estimation under different measurement noise setups (the covariance of the measurement noises differs in different setups). To ensure a fair comparison and reduce the impact of Monte Carlo randomness, each KF is simulated 10,000 times for each measurement-noise level. All filters share the same realizations of the random initial state errors, process noises, and measurement noises. Concretely, these random sequences are generated from the same random seed so that the $i^{\text{th}}$ simulation of each method (e.g., EKF, UKF, CKF) is driven by identical noise realizations.
@@
3.1 State estimation accuracy
The root mean squared error (RMSE) of the state estimations given by different nonlinear KFs are shown in Figs. 37. In all figures, different types of nonlinear KFs are represented by different colors and markers, and the old and new frameworks are represented by dotted lines and solid lines, respectively. The x-axis represents the standard deviations of the measurements. The root mean squared error is calculated from the estimation errors in 10,000 simulations. The estimation error is calculated as the difference between the estimated state and the true state at the end of the final iteration of each simulation. Note that we only show the RMSE of two representative states for systems with more than two states. For example, in 3D target tracking, the six states are the object’s speed and position along the three axes, and Fig. 3 examines the object’s speed and position estimation accuracy along the x-axis. As can be seen from the five figures, the simulation results are very similar to each other. Therefore, we mainly analyze the first application, 3D target tracking, in depth in this section.
@@ 3.1 State estimation accuracy
Figure 5: The root mean squared error of the state estimations under different measurement noise setups (synchronous generator state estimation).
@@
By comparing the performance of IEKF and EKF (with and without the proposed framework) in Figs. 37, we observe that although the IEKF outperforms the EKF under the conventional framework, it is consistently inferior to the EKF implemented with the new framework, while also incurring higher computational cost. As discussed in the Introduction, this discrepancy arises because the IEKF continues to produce overconfident covariance estimates, particularly in cases where the iteration fails to converge — a situation that occurs frequently in practice.
The necessity of the “back out” step is illustrated in Fig. 8, which compares the RMSEs of pendulum state estimation (Application 3) with and without this step. The results show that, without the “back out” step, all four nonlinear KF variants become significantly less accurate when the measurement noise is low. This degradation can be attributed to the fact that the a posteriori covariance computed in (25) is not guaranteed to be bounded. As illustrated previously in Fig. 2, without the “back out” step, the covariance may grow explosively, leading to severely inaccurate state estimates in subsequent iterations. It should be noted, however, that the impact of the “back out” step is application dependent. For example, as shown in Fig. 9, removing this step in Application 1 noticeably degrades only the EKF performance, while EKF2, UKF, and CKF are almost unaffected. Nevertheless, it is still advisable to retain this step to guard against potential covariance explosion, since it incurs negligible computational overhead in the proposed framework. Appendix D presents the corresponding results for the other applications and reports, for all five applications, the percentage of update steps in which the “back out” mechanism is actually triggered.
Figure 8: The root mean squared error of the state estimations with and without the “back out” step under different measurement noise setups (pendulum state estimation).
Figure 9: The root mean squared error of the state estimations with and without the “back out” step under different measurement noise setups (3D target tracking).
3.2 Covariance estimation accuracy
As discussed previously, the proposed framework aims to improve covariance estimation by re-approximating the measurement functions after each state update. To evaluate whether the framework actually leads to more accurate covariance estimates, we introduce two metrics for assessing covariance estimation accuracy: the Average Normalized Estimation Error Squared (ANEES) and the Noncredibility Index (NCI). Specifically, the ANEES at time step $k$ is defined as [35]:
$\mathrm{ANEES}_{k}:=\frac{1}{n_{x}N_{\text{MC}}}\sum_{i=1}^{N_{\text{MC}}}\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)^{T}\bigl({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}}\bigr)^{-1}\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr),$ (33)
where $n_{x}$ is the state dimension and $N_{\text{MC}}=10{,}000$ is the number of Monte Carlo simulations. The superscript $(i)$ denotes the $i^{\text{th}}$ Monte Carlo simulation. Here, $x_{k}^{(i)}$ and $\hat{x}_{k|k}^{(i)}$ are the true and estimated states at time step $k$, respectively, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}$} is the corresponding state estimation error covariance matrix. Note that ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k|k-1,\mathrm{est}}}$ for the conventional framework, whereas ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k|k,\mathrm{est}}}$ for the proposed framework. When $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}$} equals the true error covariance and the estimation error is zero-mean Gaussian, we have $\operatorname{\mathbb{E}}[\mathrm{ANEES}_{k}]=1$. Therefore, the deviation of $\mathrm{ANEES}_{k}$ from 1 reflects the degree of overconfidence or conservatism in the KF. However, since the estimation error can be biased and non-Gaussian in practice, the value of ANEES alone may not be a sufficiently reliable indicator of overconfidence [35]. Therefore, we also consider the metric NCI, which quantifies the mismatch between the estimated covariance and the empirical covariance of the estimation error. Specifically, the NCI at time step $k$ is defined as [36]:
$\mathrm{NCI}_{k}:=\frac{10}{N_{\text{MC}}}\sum_{i=1}^{N_{\text{MC}}}\log_{10}\frac{\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)^{T}\bigl({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}}\bigr)^{-1}\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)}{\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)^{T}\bigl(P^{*}_{k}\bigr)^{-1}\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)},$ (34)
where $P^{*}_{k}$ is the sample error covariance at time step $k$, defined by
$P^{*}_{k}:=\frac{1}{N_{\text{MC}}}\sum_{i=1}^{N_{\text{MC}}}\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)^{T}.$ (35)
By construction, $\mathrm{NCI}_{k}\approx 0$ when $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}$} matches $P_{k}^{*}$, $\mathrm{NCI}_{k}>0$ indicates overconfidence, and $\mathrm{NCI}_{k}<0$ indicates conservatism. Given the definitions of ANEES and NCI for each step $k$, we define the global ANEES and NCI as the average across all time steps:
$\begin{cases}\mathrm{ANEES}=\dfrac{1}{N_{\text{step}}}\displaystyle\sum_{k=1}^{N_{\text{step}}}\mathrm{ANEES}_{k},\\[8.00003pt] \mathrm{NCI}=\dfrac{1}{N_{\text{step}}}\displaystyle\sum_{k=1}^{N_{\text{step}}}\mathrm{NCI}_{k},\end{cases}$ (36)
where $N_{\text{step}}$ is the total number of time steps in the simulation. Given these metrics, Fig. 10 compares the global ANEES and NCI of various KF algorithms using the conventional and proposed frameworks in the target tracking application. Under the conventional framework, when the standard deviations of both measurements are below 0.1 m, all nonlinear KFs substantially underestimate the covariance, regardless of the specific algorithm. In contrast, under the proposed framework, only the EKF remains noticeably overconfident, while the ANEES and NCI of the other nonlinear KFs are very close to one and zero, respectively. This residual gap for the EKF is consistent with its neglect of second- and higher-order terms in the measurement functions; nevertheless, its covariance estimates are still significantly more accurate when combined with the proposed framework.
Figure 10: The global Average Normalized Estimation Error Squared (ANEES) and Noncredibility Index (NCI) of various nonlinear Kalman filters under different measurement noise setups. The new framework significantly mitigates the problem of overconfidence.
In Fig. 11, we compare the estimated standard deviation of the first state $\hat{\sigma}_{1}$ using different KF variants and frameworks in the target tracking application. Specifically, $\hat{\sigma}_{1}$ is defined as the estimated posterior standard deviation of the $x$-axis position in the final iteration, averaged over all Monte Carlo simulations. Mathematically,
$\hat{\sigma}_{1}:=\frac{1}{N_{\text{MC}}}\sum_{i=1}^{N_{\text{MC}}}\sqrt{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{N_{\text{step}},\mathrm{est}}}(1,1)}.$ (37)
Figure 11: Estimated standard deviation of the first state for various nonlinear Kalman filters under different measurement noise setups.
Interestingly, Fig. 11 suggests that the estimated standard deviation is almost the same for both frameworks, regardless of the KF algorithm. These results support the conclusion discussed in Section 2.4: the key to achieving more accurate state estimation is to reduce the gap between the estimated covariance and the actual covariance, rather than merely finding a better Kalman gain that minimizes the estimated covariance.
3.3 Convergence characteristics
@@
Another important metric for evaluating the performance of different filters is the convergence characteristics. We visualize this metric by fixing the standard deviations of the two measurements to 0.01 m and plotted the RMSE of the state estimations after different numbers of iterations. The results are shown in Fig. 12. From the figure, we can see that all types of nonlinear KFs can converge faster after being combined with the new framework. This result suggests that the proposed framework is very suitable for applications that require high convergence speed.
@@
3.4 Computational time
Since the extra steps introduced in the new framework take time, it is also important to compare the computational time of the two frameworks. The results are summarized in Fig. 13. Note that the average runtime shown in the figure is calculated from over 100,000 simulations run on the same computer. Specifically, for each combination of nonlinear KF and framework, we record the method’s runtime whenever it was used and calculated the average. After the average runtime is calculated, the value is normalized with respect to the average runtime of the conventional EKF for the same application. In other words, the normalized average runtime of “EKF (old)” always equals one in Fig. 13.
Figure 13: Comparison of the normalized computational time of different methods and frameworks.
@@ 4 Conclusion
In this paper, we propose a new framework that can be applied to all types of nonlinear KFs. The new framework re-approximates the measurement functions after the state update, solving the problem of overconfident state covariance estimation, which is an inherent problem in the conventional framework for nonlinear KFs. The new framework was tested in five different applications, reducing the estimation errors by more than an order of magnitude when the measurement noise is low and significantly reducing the convergence time of the algorithm. Considering that all types of sensors will become increasingly more accurate in the future, the proposed framework is expected to bring greater benefits for all kinds of nonlinear KFs over time.
@@ Appendix A Supplementary proof for Theorem 1
[…]
@@
Positive-definiteness is satisfied because for any $x\in L^{2}$, $\mathbb{E}[x^{T}x]\geq 0$, and the equality sign only holds when $x=0$ almost surely.
@@
Let $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S$} be a random symmetric and positive definite matrix whose mean is the identity matrix $I$. Let $\bm{v}$ be a random vector whose mean is $\bm{v}_{0}$. The following inequality holds:
$\mathbb{E}(\bm{v}^{T}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}\bm{v})\geq\bm{v}_{0}^{T}\bm{v}_{0}.$ (38)
Note that the equality sign only holds when ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}\bm{v}$ is constant almost surely.
@@
Let $\bm{x}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-\frac{1}{2}}}\bm{v},\bm{y}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{\frac{1}{2}}}\bm{v}_{0}$. Since $\bm{v}$ is a random vector with a mean of $\bm{v}_{0}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S$} is a random symmetric matrix with a mean of the identity matrix,
$\begin{cases}\mathbb{E}[\bm{x}^{T}\bm{x}]=\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}\bm{v}]\\ \mathbb{E}[\bm{y}^{T}\bm{y}]=\mathbb{E}[\bm{v}_{0}^{T}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S}\bm{v}_{0}]=\bm{v}_{0}^{T}\mathbb{E}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S}]\bm{v}_{0}=\bm{v}_{0}^{T}\bm{v}_{0}\\ \mathbb{E}[\bm{x}^{T}\bm{y}]=\mathbb{E}[\bm{v}^{T}\bm{v}_{0}]=\mathbb{E}[\bm{v}]^{T}\bm{v}_{0}=\bm{v}_{0}^{T}\bm{v}_{0}\end{cases}$ (39)
According to Lemma 3, $\mathbb{E}[\bm{x}^{T}\bm{y}]$ is an inner product on $L^{2}(\Omega;\mathbb{R}^{n})$. Applying Cauchy–Schwarz inequality, we have
$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}\bm{v}]\geq\frac{(\bm{v}_{0}^{T}\bm{v}_{0})^{2}}{\bm{v}_{0}^{T}\bm{v}_{0}},$ (40)
which completes the proof. Note that the equality holds only when ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}\bm{v}$ is constant almost surely.
@@ Appendix B The detailed simulation setup
This example is adapted from Example 3.3.1 in [38]. As shown in Fig. 14(a), a target moves in 3D space, and two sensors are used to track its location. The first sensor is at the origin, and the second sensor moves in a circle. The second sensor’s location $(x_{s1},x_{s2},x_{s3})$ at step $k=1,2,\ldots,30$ is represented by
$\begin{cases}x_{s1,k}=20+20\cos{\big(\frac{(k-1)\pi}{15}\big)},\\ x_{s2,k}=20+20\sin{\big(\frac{(k-1)\pi}{15}\big)},\\ x_{s3,k}=0.\end{cases}$ (41)
The states are the target’s positions $(x_{1},x_{2},x_{3})$ and speeds $(v_{1},v_{2},v_{3})$ along the three axes. The measurements $z_{1},z_{2}$ are the distances between the target and the two sensors. The inputs are the target’s accelerations along the three axes, which are assumed to be zero in this example. When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (42) and (43). Note that the noises are omitted in the equations.
$\begin{bmatrix}x_{1,k}\\ x_{2,k}\\ x_{3,k}\\ v_{1,k}\\ v_{2,k}\\ v_{3,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}+v_{1,k-1}\\ x_{2,k-1}+v_{2,k-1}\\ x_{3,k-1}+v_{3,k-1}\\ v_{1,k-1}\\ v_{2,k-1}\\ v_{3,k-1}\end{bmatrix}$, (42)
$\begin{bmatrix}z_{1,k}\\ z_{2,k}\end{bmatrix}=\begin{bmatrix}\sqrt{\sum_{i=1}^{3}x_{i,k}^{2}}\\ \sqrt{\sum_{i=1}^{3}(x_{i,k}-x_{si,k})^{2}}\end{bmatrix}$. (43)